1-5 The ball is tied to a string to whirl it having a radius of 40.0 cm at a velocity of 2.5 m/s. What will be




the acceleration of the ball?

1
Expert's answer
2022-03-25T15:02:22-0400

The centripetal acceleration 

"a=\frac{v^2}{R}=\frac{2.5^2}{0.20}=31\:\rm m/s^2"